Calculate the grams of water that must be added to 16.0 grams of sugar in the preparation of a 25.0% sugar solution

1
Expert's answer
2021-11-26T06:34:02-0500

"%=\frac{m_{sugar}{" %"=\frac{m_{(sugar)}}{m_{(sugar)}+m(H_2O)}×100"


25%"=\frac{16}{16+m(H_2O)}×100"


"m(H_2O)=\frac{16×100}{25}-16=48gram"